OpenStudy (anonymous):

-x^3 -125 factor completely.

OpenStudy (tyteen4a03):

Find the cube root of 125, then apply the cube identity: \(x^3 - y^3 = (x-y)(x^2+xy+y^2)\).

OpenStudy (tyteen4a03):

(Note, the cube identity has another version: \(x^3 + y^3 = (x+y)(x^2-xy+y^2)\))